# Quantum many-body scars with unconventional superconducting pairing symmetries via multi-body interactions

### v2

@@ III.1 Spinless $s$-wave pairing state in a one-dimensional system
with the interaction strength $W$. We observe that when two pairs come next to each other there always appear two combinations of neighboring three particles [Fig. 1(b)]. This motivates us to choose the three-body interaction parameter $W=-V/2$, which cancels the energy increase due to the two-body interaction. The cancellation works for arbitrary number $N/2$ of pairs no matter how they distribute on the lattice. It also works for both the periodic and open boundary conditions. In the latter case, the interaction terms should be defined as $\mathcal{H}_{V}=V\sum_{i=1}^{L-1}n_{i}n_{i+1}$ and $\mathcal{H}_{W}=\sum_{i=1}^{L-2}n_{i}n_{i+1}n_{i+2}$ with $L$ being the number of lattice sites. Hence, the spinless $s$-wave $\eta$-pairing state becomes an exact eigenstate of the Hubbard model with the three-body interaction (with $W=-V/2$) under the periodic or open boundary condition,
@@ III.3 Hubbard model with the two-body interaction
$\mathcal{H}_{\mathrm{kin}}=-t\sum_{\langle i,j\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma},$ (21)
@@
The $d$-wave $\eta$-pairing state is expanded as a linear combination of Fock states in which various oriented pairs are distributed on the square lattice. Examples of those Fock states are shown in Fig. 2. When two off-site pairs are present, they are either un-overlapped or overlapped, as shown in Fig. 2(a) and (b), respectively. In the latter case, a doublon is formed, generating an energy cost due to the on-site two-body interaction. This means that the latter Fock basis [Fig. 2(b)] has higher interaction energy (with the difference $U$) than the former basis [Fig. 2(a)]. Therefore, the $d$-wave $\eta$-pairing state $|\Psi_{d}^{N}\rangle$ in Eq. (19) is not an energy eigenstate of $\mathcal{H}_{U}$ [Eq. (22)]. When three (or more) pairs exist, more than two particles can occupy nearest-neighbor sites of a doublon, as shown in Fig. 2(c). This example indicates that a naive application of the three-body interaction introduced in Sec. III.1 is not sufficient to cancel the energy increment of the $d$-wave pairs due to the formation of doublons. The question is how such an energy change can be cancelled for all possible configurations of off-site pairs.
III.4 Hubbard model with multi-body interactions
Here we show that the unconventional $\eta$-pairing states can become exact energy eigenstates by introducing generalized multi-body interactions to the Hubbard model. As mentioned before, our motivation is to cancel the energy increment due to the formation of doublons by using the multi-body interactions. Since the number of particles that surround a doublon may vary depending on the configurations of pairs, but at least two are present (see Fig. 2), it will be convenient to introduce an operator that judges whether neighboring sites (next to a doublon) are occupied by particles or not. This allows us to efficiently measure how much energy should be subtracted for each pair configuration to make the unconventional $\eta$-pairing states exact eigenstates.
@@ III.4 Hubbard model with multi-body interactions
Figure 4: Interactions between off-site pairs in the $d$-wave $\eta$-pairing state in a spinful model with spin-$\uparrow$ (spin-$\downarrow$) particles indicated by the red (blue) arrows. (a) When two pairs are overlapped with each other to form a doublon, the on-site two-body interaction is cancelled by the three-body interaction [Eq. (27)] acting between the doublon and a single particle. (b) When another pair comes next to the doublon, the on-site two-body interaction is not cancelled by the three-body interaction [Eq. (27)], but is cancelled by the multi-body interaction in the form of Eq. (28). Two particles forming a doublon are indicated by the red squares, while the multi-body interacting doublon and spin-$\uparrow$ (spin-$\downarrow$) particles are enclosed by the blue (green) boxes.
In order to get an insight of how to cancel the energy increment of doublons with multi-body interactions, we take two examples of Fock states relevant to the unconventional $\eta$-pairing states: One is the case where two pairs are overlapped to each other to form a doublon, as shown in Fig. 4(a). The other is the case where there is another pair in the vicinity of the two pairs, as shown in Fig. 4(b). We label the site occupied by the doublon by $i$. As shown in Fig. 4(a), there are always two residual particles from the pairs that form the doublon, so that we can cancel the on-site two-body interaction $Un_{i\uparrow}n_{i\downarrow}$ by adding the following three-body interaction between the doublon and a particle,
@@
The multi-body interaction of Eq. (28) contains $M$-body interactions with $23\leq M\leq z+$ 2. With this form of the multi-body interaction, it is possible to incorporate only the interaction between the doublon and a single spin-$\uparrow$ particle, and the interaction between the doublon and a single spin-$\downarrow$ particle. One can see that the cancellation between the on-site two-body interaction $\mathcal{H}_{U}$ [Eq. (22)] and the multi-body interaction [Eq. (28)] works not only for the cases of Fig. 4(a) and (b) but also for all the possible configurations of pairs in the unconventional $\eta$-pairing states, which can be easily checked since the cancellation occurs for each doublon one by one.
Based on the multi-body interaction in Eq. (28) introduced above, we define an extended Hubbard model having the unconventional $\eta$-pairing states as energy eigenstates. The Hamiltonian of the extended Hubbard model, denoted by $\mathcal{H}_{\mathrm{extH}}$, is given as follows:
@@
As discussed in the previous paragraph, the interaction energy is always zero for the unconventional $\eta$-pairing states, due to the cancellation between the two-body interaction and the multi-body interaction [Eq. (28)]. In particular, the $d$-wave $\eta$-pairing state $|\Psi_{d}^{N}\rangle$ in Eq. (19) is an exact eigenstate of $\mathcal{H}_{\mathrm{extH}}$ with a zero eigenenergy,
@@ IV.2 Entanglement entropy
Next, we verify a nonthermal behavior of the $d$-wave $\eta$-pairing state through the entanglement entropy. We consider the same $L_{x}\times L_{y}$ lattice as in the previous subsection with the periodic boundary condition. The entire system (a torus) is divided into subsystems $A$ and $B$ (two tubes), where $A$ has the size of $r_{x}\times L_{y}$. The entanglement entropy for the $d$-wave $\eta$-pairing state in Eq. (19) is defined by $S_{A}(|\Psi_{d}^{N}\rangle)=-\mathrm{Tr}_{A}[\rho_{A}\ln\rho_{A}]$, where $\rho_{A}$ is the reduced density matrix for the subsystem $A$, i.e., $\rho_{A}={\mathrm{Tr}}_{B}[|\Psi_{d}^{N}\rangle\langle\Psi_{d}^{N}|]$. We compute $S_{A}(|\Psi_{N,d}\rangle)$ as a function of the volume fraction $f=V_{A}/V$, where $V=2L_{x}L_{y}$ ($V_{A}=2r_{x}L_{y}$) is the product of the system (subsystem) volume and the spin degrees of freedom.
@@
Figure 6 shows the calculated entanglement entropy for the $d$-wave $\eta$-pairing state indicated by the blue cross points. We also show the average entanglement entropy $\langle S_{A}\rangle_{N}$ and $\langle S_{A}\rangle_{G,N}$ with red dashed and green chained curves, respectively. We observe that the entanglement entropy of the $d$-wave $\eta$-pairing state increases as the subsystem volume increases. Particularly, we found that the subsystem size dependence of the entanglement entropy is in quantitative agreement with that typically observed in integrable systems, which is substantially smaller than the typical entanglement entropy in nonintegrable systems. These numerical results support the nonthermal property of the $d$-wave $\eta$-pairing state.
V Nonintegrability of the multi-body interacting model
@@
In this section, we numerically confirm the nonintegrability of the extended Hubbard model $\mathcal{H}_{\mathrm{extH}}$ in Eq. (29) from the energy level statistics. Combining the results with those of Sec. IV, we can judge whether the $d$-wave $\eta$-pairing state in the model is identified to be a QMBS state or not. If the distribution of the difference between the nearest-neighbor eigenenergies follows the Wigner-Dyson distribution $P_{\mathrm{WD}}(s)=(\pi/2)s\exp(-\pi s^{2}/4)$ with $s$ being the level spacing, then the model is suggested to be nonintegrable and all the other energy eigenstates are thermal as expected from the argument of the ETH.
@@
We consider the extended Hubbard model in Eq. (29) on the square lattice under the periodic boundary condition with $L_{x}=L_{y}=4$, $N=8$, and the total magnetization $S^{z}=0$. We utilize the Lanczos method to numerically diagonalize the Hamiltonian. Since the Hamiltonian has several global symmetries, we select a target subspace using symmetry projection operators. As internal symmetries, we consider the time-reversal ($\mathcal{T}$) and spin-rotation symmetries. It should be noted that our multi-body interacting system does not have the particle-hole symmetry, unlike the two-body interacting Hubbard model in Eq. (20). Among the elements of the point group $C_{4v}$ (which is the symmetry of the square lattice), we take the reflections $\mathcal{P}_{x}$, $\mathcal{P}_{y}$, and $\mathcal{P}_{d}$ with respect to the $x$-, $y$-, and diagonal axes, respectively. The fixed point of the point group operations is taken to be a site center, respecting the staggered phase oscillation of the $\eta$-pairing state. Additionally, the system has the translation symmetries $\mathcal{X}$ and $\mathcal{Y}$ along the $x$ and $y$ directions, respectively. We focus on the subspaces in which the $d$-wave $\eta$-pairing state exists, namely with $\mathcal{T}=+1$, the total spin $S=0$, $(\mathcal{P}_{x},\mathcal{P}_{y},\mathcal{P}_{d})=(+1,+1,+1)$, and the total momentum $(0,0)$. The last three are for the cases where $N/2$ is even. The $d$-wave $\eta$-pairing state belongs to the total-spin eigenspace with $S=0,2,4$ when $N=8$. Note that in a $4\times 4$ periodic lattice system, there is an accidental hidden symmetry, i.e., the four-dimensional hypercubic symmetry [77]. If this higher symmetry is not taken into account to define the target space, the level-spacing distributions may behave differently from the Wigner-Dyson distribution (even though the system might be nonintegrable), which will be discussed later.
@@
Figure 7: Nearest-neighbor level-spacing distribution $P(s)$ (shown by the histograms) for the extended Hubbard model with the multi-body interactions [Eq. (29)] on the square lattice with the quantum numbers (a) $(\mathcal{P}_{x},\mathcal{P}_{y},\mathcal{P}_{d})=(+1,+1,+1)$ and (b) $(+1,+1,-1)$. The parameters are set to be $L_{x}=L_{y}=4$, $N=8$, and $U/t=2$. The blue and red curves correspond to the Poisson and Wigner-Dyson distributions, respectively.
@@
Figure 7(a) shows the nearest-neighbor level-spacing distribution $P(s)$ in the subspace with the quantum numbers $(\mathcal{P}_{x},\mathcal{P}_{y},\mathcal{P}_{d})=(+1,+1,+1)$, where we adopt the unfolding method with the Gaussian kernel density estimation [78] with the smoothing parameter $\sigma_{\mathrm{Gauss}}=0.1$. We find that the obtained level-spacing distribution has a shape in between the Poisson distribution $P_{\mathrm{P}}(s)=\exp(-s)$ and the Wigner–Dyson distribution $P_{\mathrm{WD}}(s)$. A similar behavior is often observed when an extra discrete symmetry remains to be considered in a nonintegrable model. For example, in the two-body interacting Hubbard model of Eq. (20) on the same $4\times 4$ periodic lattice, the level-spacing distribution has been found to be neither the Poisson nor Wigner-Dyson distribution [77]. As we remarked before, this is due to the presence of the accidental hypercubic symmetry for the $4\times 4$ lattice geometry, which has not been taken into account in the level-spacing analysis. Such an intermediate distribution between the Poisson and Wigner-Dyson ones has been reproduced by a spectrum of two independent GOE samples in the random matrix theory mixed with appropriate weights [77]. We also examine the subspace with another set of the quantum numbers $(\mathcal{P}_{x},\mathcal{P}_{y},\mathcal{P}_{d})=(+1,+1,-1)$, and obtain the level-spacing distribution as shown in Fig. 7(b). We observe that the spectrum is in good agreement with the Wigner-Dyson distribution. Based on these results, we conclude that the multi-body interacting model of Eq. (29) is nonintegrable.
@@
Given the above results with the nonthermal properties of the eigenstate seen in Sec. IV, the $d$-wave $\eta$-pairing state [Eq. (19)] can be regarded as a QMBS state in the nonintegrable system [Eq. (29)]. The basic principle that makes the unconventional $\eta$-pairing states QMBSs is to use multi-body interactions to protect the pairs from having the energy increase due to the two-body interaction.

### v3

@@ III.1 Spinless $s$-wave pairing state in a one-dimensional system
with the interaction strength $W$. We observe that when two pairs come next to each other there always appear two combinations of neighboring three particles [Fig. 1(b)]. This motivates us to choose the three-body interaction parameter $W=-V/2$, which cancels the energy increase due to the two-body interaction. The cancellation works for arbitrary number $N/2$ of pairs no matter how they distribute on the lattice. It also works for both the periodic and open boundary conditions. In the latter case, the interaction terms should be defined as $\mathcal{H}_{V}=V\sum_{i=1}^{L-1}n_{i}n_{i+1}$ and $\mathcal{H}_{W}=W\sum_{i=1}^{L-2}n_{i}n_{i+1}n_{i+2}$ with $L$ being the number of lattice sites. Hence, the spinless $s$-wave $\eta$-pairing state becomes an exact eigenstate of the Hubbard model with the three-body interaction (with $W=-V/2$) under the periodic or open boundary condition,
@@ III.3 Hubbard model with the two-body interaction
$\mathcal{H}_{\mathrm{kin}}=-t\sum_{\langle i,j\rangle\sigma}(c_{i\sigma}^{\dagger}c_{j\sigma}+\mathrm{H.c.}),$ (21)
@@
The $d$-wave $\eta$-pairing state is expanded as a linear combination of Fock states in which various oriented pairs are distributed on the square lattice. Examples of those Fock states are shown in Fig. 2. When two off-site pairs are present, they are either un-overlapped or overlapped, as shown in Fig. 2(a) and (b), respectively. In the latter case, a doublon is formed, generating an energy cost due to the on-site two-body interaction. This means that the latter Fock basis [Fig. 2(b)] has higher interaction energy (with the difference $U$) than the former basis [Fig. 2(a)]. Therefore the $d$-wave $\eta$-pairing state $|\Psi_{d}^{N}\rangle$ in Eq. (19) is not an energy eigenstate of $\mathcal{H}_{U}$ [Eq. (22)]. When three (or more) pairs exist, more than two particles can occupy nearest-neighbor sites of a doublon, as shown in Fig. 2(c). This example indicates that a naive application of the three-body interaction introduced in Sec. III.1 is not sufficient to cancel the energy increment of the $d$-wave pairs due to the formation of doublons. The question is how such an energy change can be canceled for all possible configurations of off-site pairs.
III.4 Hubbard model with multibody interactions
Here we show that the unconventional $\eta$-pairing states can become exact energy eigenstates by introducing generalized multibody interactions to the Hubbard model. As mentioned before, our motivation is to cancel the energy increment due to the formation of doublons by using the multibody interactions. Since the number of particles that surround a doublon may vary depending on the configurations of pairs, but at least two are present (see Fig. 2), it will be convenient to introduce an operator that judges whether neighboring sites (next to a doublon) are occupied by particles or not. This allows us to efficiently measure how much energy should be subtracted for each pair configuration to make the unconventional $\eta$-pairing states exact eigenstates.
@@ III.4 Hubbard model with multibody interactions
Figure 4: Interactions between off-site pairs in the $d$-wave $\eta$-pairing state in a spinful model with spin-$\uparrow$ (spin-$\downarrow$) particles indicated by the red (blue) arrows. (a) When two pairs are overlapped with each other to form a doublon, the on-site two-body interaction is canceled by the three-body interaction [Eq. (27)] acting between the doublon and a single particle. (b) When another pair comes next to the doublon, the on-site two-body interaction is not canceled by the three-body interaction [Eq. (27)], but is canceled by the multibody interaction in the form of Eq. (28). Two particles forming a doublon are indicated by the red squares, while the multibody interacting doublon and spin-$\uparrow$ (spin-$\downarrow$) particles are enclosed by the blue (green) boxes.
In order to get an insight of how to cancel the energy increment of doublons with multibody interactions, we take two examples of Fock states relevant to the unconventional $\eta$-pairing states: One is the case where two pairs are overlapped to each other to form a doublon, as shown in Fig. 4(a). The other is the case where there is another pair in the vicinity of the two pairs, as shown in Fig. 4(b). We label the site occupied by the doublon by $i$. As shown in Fig. 4(a), there are always two residual particles from the pairs that form the doublon, so that we can cancel the on-site two-body interaction $Un_{i\uparrow}n_{i\downarrow}$ by adding the following three-body interaction between the doublon and a particle,
@@
The multibody interaction of Eq. (28) contains $M$-body interactions with $23\leq M\leq z+$ 2. With this form of the multibody interaction, it is possible to incorporate only the interaction between the doublon and a single spin-$\uparrow$ particle, and the interaction between the doublon and a single spin-$\downarrow$ particle. One can see that the cancellation between the on-site two-body interaction $\mathcal{H}_{U}$ [Eq. (22)] and the multibody interaction [Eq. (28)] works not only for the cases of Fig. 4(a) and (b) but also for all the possible configurations of pairs in the unconventional $\eta$-pairing states, which can be easily checked since the cancellation occurs for each doublon one by one.
Based on the multibody interaction in Eq. (28) introduced above, we define an extended Hubbard model having the unconventional $\eta$-pairing states as energy eigenstates. The Hamiltonian of the extended Hubbard model, denoted by $\mathcal{H}_{\mathrm{extH}}$, is given as follows:
@@
As discussed in the previous paragraph, the interaction energy is always zero for the unconventional $\eta$-pairing states, due to the cancellation between the two-body interaction and the multibody interaction [Eq. (28)]. In particular, the $d$-wave $\eta$-pairing state $|\Psi_{d}^{N}\rangle$ in Eq. (19) is an exact eigenstate of $\mathcal{H}_{\mathrm{extH}}$ with a zero eigenenergy,
@@
We comment on the robustness against perturbations on the parameters of the multibody interactions. When the two off-site pairs coexist, as shown in Fig. 4(a), the three-body interaction in Eq. (27) is sufficient to cancel the on-site two-body interaction $U$ associated with the doublon formation. However, when an additional pair comes next to this doublon, as shown in Fig. 4(b), higher multibody interactions in Eq. (28) become active. These higher-order terms are generally required only when off-site pairs are densely distributed. In cases with a low pair density, off-site pairs do not frequently come close to doublons, and higher multibody interactions become practically irrelevant. Thus, we do not have to fine-tune the parameters of the higher multibody interactions.
@@ IV.2 Entanglement entropy
Next, we verify a nonthermal behavior of the $d$-wave $\eta$-pairing state through the entanglement entropy. We consider the same $L_{x}\times L_{y}$ lattice as in the previous subsection with the periodic boundary condition. The entire system (a torus) is divided into subsystems $A$ and $B$ (two tubes), where $A$ has the size of $r_{x}\times L_{y}$. The entanglement entropy for the $d$-wave $\eta$-pairing state in Eq. (19) is defined by $S_{A}(|\Psi_{d}^{N}\rangle)=-\mathrm{Tr}_{A}[\rho_{A}\ln\rho_{A}]$, where $\rho_{A}$ is the reduced density matrix for the subsystem $A$, i.e., $\rho_{A}={\mathrm{Tr}}_{B}[|\Psi_{d}^{N}\rangle\langle\Psi_{d}^{N}|]$. We compute $S_{A}(|\Psi_{d}^{N}\rangle)$ as a function of the volume fraction $f=V_{A}/V$, where $V=2L_{x}L_{y}$ ($V_{A}=2r_{x}L_{y}$) is the product of the system (subsystem) volume and the spin degrees of freedom.
@@
Figure 6 shows the calculated entanglement entropy for the $d$-wave $\eta$-pairing state indicated by the blue cross points. We also show the average entanglement entropy $\langle S_{A}\rangle_{N}$ and $\langle S_{A}\rangle_{G,N}$ with red dashed and green chained curves, respectively. We observe that the entanglement entropy of the $d$-wave $\eta$-pairing state increases as the subsystem volume increases. Particularly, we found that the subsystem size dependence of the entanglement entropy is in quantitative agreement with that typically observed in integrable systems, which is substantially smaller than the typical entanglement entropy in nonintegrable systems. These numerical results support the nonthermal property of the $d$-wave $\eta$-pairing state. In this study, due to the constraint on a computational cost, we have not examined the volume dependence of the entanglement entropy in the thermodynamic limit. However, similarly to the Yang’s $\eta$-pairing state [52] and the spinless $\eta$-pairing state [64], we can expect that the entanglement entropy of the $d$-wave $\eta$-pairing state follows a logarithmic dependence on the subvolume.
V Nonintegrability of the multibody interacting model
@@
In this section, we numerically confirm the nonintegrability of the extended Hubbard model $\mathcal{H}_{\mathrm{extH}}$ in Eq. (29) from the energy level statistics. Combining the results with those of Sec. IV, we can judge whether the $d$-wave $\eta$-pairing state in the model is identified to be a QMBS state or not. If the distribution of the difference between the nearest-neighbor eigenenergies follows the Wigner–Dyson distribution $P_{\mathrm{WD}}(s)=(\pi/2)s\exp(-\pi s^{2}/4)$ with $s$ being the level spacing, then the model is suggested to be nonintegrable and all the other energy eigenstates are thermal as expected from the argument of the ETH.
@@
We consider the extended Hubbard model in Eq. (29) on the square lattice under the periodic boundary condition with $L_{x}=L_{y}=4$, $N=8$, and the total magnetization $S^{z}=0$. We utilize the Lanczos method to numerically diagonalize the Hamiltonian. Since the Hamiltonian has several global symmetries, we select a target subspace using symmetry projection operators. As internal symmetries, we consider the time-reversal ($\mathcal{T}$) and spin-rotation symmetries. It should be noted that our multibody interacting system does not have the particle-hole symmetry, unlike the two-body interacting Hubbard model in Eq. (20). Among the elements of the point group $C_{4v}$ (which is the symmetry of the square lattice), we take the reflections $\mathcal{P}_{x}$, $\mathcal{P}_{y}$, and $\mathcal{P}_{d}$ with respect to the $x$-, $y$-, and diagonal axes, respectively. The fixed point of the point group operations is taken to be a site center, respecting the staggered phase oscillation of the $\eta$-pairing state. Additionally, the system has the translation symmetries $\mathcal{X}$ and $\mathcal{Y}$ along the $x$ and $y$ directions, respectively. We focus on the subspaces in which the $d$-wave $\eta$-pairing state exists, namely with $\mathcal{T}=+1$, the total spin $S=0$, $(\mathcal{P}_{x},\mathcal{P}_{y},\mathcal{P}_{d})=(+1,+1,+1)$, and the total momentum $(0,0)$. The last three are for the cases where $N/2$ is even. The $d$-wave $\eta$-pairing state belongs to the total-spin eigenspace with $S=0,2,4$ when $N=8$. Note that in a $4\times 4$ periodic lattice system, there is an accidental hidden symmetry, i.e., the four-dimensional hypercubic symmetry [78]. If this higher symmetry is not taken into account to define the target space, the level-spacing distributions may behave differently from the Wigner–Dyson distribution (even though the system might be nonintegrable), which will be discussed later.
@@
Figure 7: (a) Nearest-neighbor level-spacing distribution $P(s)$ and (b) higher-order spacing ratio distribution $P^{k=2}(r)$ (shown by the histograms) for the extended Hubbard model with the multibody interactions [Eq. (29)] on the square lattice with the quantum numbers $(\mathcal{P}_{x},\mathcal{P}_{y},\mathcal{P}_{d})=(+1,+1,+1)$. The parameters are set to be $L_{x}=L_{y}=4$, $N=8$, and $U/t=2$.
@@
Figure 7(a) shows the nearest-neighbor level-spacing distribution $P(s)$ in the subspace with the quantum numbers $(\mathcal{P}_{x},\mathcal{P}_{y},\mathcal{P}_{d})=(+1,+1,+1)$, where we adopt the unfolding method with the Gaussian kernel density estimation [79] with the smoothing parameter $\sigma_{\mathrm{Gauss}}=0.1$. We find that the obtained level-spacing distribution has a shape in between the Poisson distribution $P_{\mathrm{P}}(s)=\exp(-s)$ and the Wigner–Dyson distribution $P_{\mathrm{WD}}(s)$. A similar behavior is often observed when an extra discrete symmetry remains to be considered in a nonintegrable model. For example, in the two-body interacting Hubbard model of Eq. (20) on the same $4\times 4$ periodic lattice, the level-spacing distribution has been found to be neither the Poisson nor Wigner–Dyson distribution [78]. As we remarked before, this is due to the presence of the accidental hypercubic symmetry for the $4\times 4$ lattice geometry, which has not been taken into account in the level-spacing analysis. Such an intermediate distribution between the Poisson and Wigner–Dyson ones has been reproduced by a spectrum of two independent GOE samples in the random matrix theory mixed with appropriate weights [78].
To extract the nonintegrability in systems with such a residual symmetry, we investigate higher-order spectral statistics [80, 81, 82]. We define the distribution $P^{k}(r)$ of the $k$th order spacing ratio with $r=(E_{i+2k}-E_{i+k})/(E_{i+k}-E_{i})$ ($E_{i}$ is the $i$th energy eigenvalue). Given superposed $m$ independent GOE samples, it has been numerically shown that $P^{k}(r)=P(r,\beta^{\prime})$ holds with the relationship $\beta^{\prime}=m=k$ [80], where $P(r,\beta)=C_{\beta}\frac{(r+r^{2})^{\beta}}{(1+r+r^{2})^{1+3\beta/2}}$ and $C_{\beta}$ is a normalization constant. In the present case of $m=2$, it is given by $C_{2}=81\sqrt{3}/4\pi$ [83]. In contrast, for integrable systems, the $k$th order spacing ratio is given by the distribution function $P^{k}_{\mathrm{P}}(r)=\frac{(2k-1)!}{[(k-1)!]^{2}}\frac{r^{k-1}}{(1+r)^{2k}}$ [80]. Figure 7(b) presents a histogram of the numerically obtained distribution of the $k$th order spacing ratio ($k=2$) for the extended Hubbard model (29), under the same conditions as in Fig. 7(a). The results confirm good agreement with $P(r,\beta^{\prime})$, which shows the nonintegrability of the multibody interacting model of Eq. (29).
@@
Given the above results with the nonthermal properties of the eigenstate seen in Sec. IV, the $d$-wave $\eta$-pairing state [Eq. (19)] can be regarded as a QMBS state in the nonintegrable system [Eq. (29)]. The basic principle that makes the unconventional $\eta$-pairing states QMBSs is to use multibody interactions to protect the pairs from having the energy increase due to the two-body interaction.